Future Value of Multiple Annuities Assume that you contribute $220 per month to a retirement plan for 20 years. Then you are able to increase the contribution to $320 per month for another 20 years. Given a 7.0 percent interest rate, what is the value of your retirement plan after 40 years?

$178,636

$177,600

$629,552

$577,459

Present Value of an Annuity Due If the present value of an ordinary, 3-year annuity is $5,900 and interest rates are 12 percent, what's the present value of the same annuity due?

$6,608

$5,900

$3,900

$4,608

Present Value of an Annuity What is the present value of a $2,000 annuity payment over 6 years if interest rates are 10 percent?

$3,543.12

$8,710.52

$1,128.95

$9,736.84

Future Value Given a 7.50 percent interest rate, compute the year 8 future value of deposits made in years 1, 2, 3, and 4 of $1,700, $1,900, $2,200, and $2,200.

rev: 09_21_2012

$11,849.07

$9,670.15

$11,272.13

$11,022.57

Future Value of an Annuity Due If the future value of an ordinary, 7-year annuity is $6,700 and interest rates are 5 percent, what's the future value of the same annuity due?

$7,169

$6,380.95

$7,034.98

$6,700
Answer #1

1.

the value of your retirement plan after 40 years

=((220*(((1+(7%/12))^(20*12)-1)/(7%/12))))*(1+(7%/12))^(20*12)+(320*(((1+(7%/12))^(20*12)-1)/(7%/12)))

=629552

the above is the answer

2.

the present value of the same annuity due

=5900*(1+12%)

=6608

3.

present value of a $2,000 annuity payment

=2000*((1-(1+10%)^(-6))/10%)

=8710.52

4.

compute the year 8 future value

=1700*(1+7.5%)^7+1900*(1+7.5%)^6+2200*(1+7.5%)^5+2200*(1+7.5%)^4

=11849.07

the above are answer

we do only first 4 questions.